OpenStudy (anonymous):

Lidia is packing her suitcase in the dark, so as to not wake her little sister. She blindly chooses 3 of her 6 sweaters. What is the probability that she chooses her 3 favorite sweaters? (Points : 3) Option A: 1/216 Option B: 1/120 Option C: 1/20 Option D: 1/18 I chose 1/120 for this one but have a feeling its wrong.

OpenStudy (automaticloveletter):

I have a feeling it's D

OpenStudy (anonymous):

Hm. Yes it would be D, 1/18. 3 out of 6. 3x6 = 18. So yes answer is D, 1/18. Hope this helps, have a wonderful day. Tag me on any more you need help with. :) P.S your adorable lol xD

OpenStudy (anonymous):

There are 3 x 2 x 1 ways of getting her favorite sweater. There are 6 x 5 x 4 ways of getting any sweater. (3 x 2 x 1) / ( 6 x 5 x 4 )

OpenStudy (anonymous):

D

OpenStudy (automaticloveletter):

3*6=18 1/18 makes sense to me

OpenStudy (anonymous):

Wait squish the way you solved it is 1/20

OpenStudy (anonymous):

There y'go.

OpenStudy (anonymous):

So which way is the right way of solving it? 1/18 or 1/20?

OpenStudy (anonymous):

It's kinda like the question that was just solved a few minutes ago about the password thing. Same principals, different numbers with more variables.

OpenStudy (anonymous):

1/18 :)

OpenStudy (anonymous):

And if you REALLY don't believe me, check out this page and look at the second question he asks. Same principals, different variables. ¯\_(ツ)_/¯ https://answers.yahoo.com/question/index?qid=20100327222446AAgJaUZ

OpenStudy (mathstudent55):

What is the probability that in the first pick she picks one of the favorite 3? 3/6 = 1/2 What is the probability that the second pick she picks another one of the favorite 3? 2/5 What is the probability that in the third pick she pick the last of the favorite 3? 1/4 P(favorite 3 out of 6 ) = 1/2 * 2*5 * 1*4 = 2/40 = 1/20

OpenStudy (automaticloveletter):

She just wants to know if it's 1/20 or 1/18

OpenStudy (anonymous):

I wanna know how to solve it:o Surprising right xD Okay thank you guys!!

OpenStudy (anonymous):

Sure thing.

OpenStudy (automaticloveletter):

Thanks @_Squish for correcting me